How much ATP is formed from aerobic metabolism of a 24:0 fatty acid?
Answer #1

24:0 fatty acid is Lignoceric acid and has 24 carbon atoms with no unsaturation present. We are to determine the number of ATP molecules that will be generated by the aerobic metabolism of this fatty acid.

The first step in fatty acid metabolism is the activation of the fatty acid by CoA resulting in the breaking of two carbons from the fatty acid in the process. Since there are a total of 24 carbon atoms, the process can repeat until the carbon chain reaches 4 carbon atoms. Therefore, a total of (24 - 4) / 2 = 20 / 2 = 10 rounds of reaction will occur thereby producing 10 molecules of Acetyl CoA. Now when the last molecule of 4 carbon atoms undergoes activation, it can be split into two molecules of Acetly CoA. Therefore, a total of 10 + 2 = 12 molecules of Acetyl CoA will be produced, and the number of rounds of activation involved will be 10 + 1 = 11 rounds. This is shown in the figure:

23 21 17 IS 3 S-Con 25 22 20 13 12 8 4 10 rounds of ß- oxidation N 10 + Scon s-lon ß - oxidation 10 요 sol toA As-cor

Also, in addition to the 12 Acetyl-CoA molecules, the beta-oxidation also produces other high-energy molecules. For each round of beta-oxidation, one FADH2 and one NADH molecules are produced. Therefore, a total of 11 FADH2 and a total of 11 NADH are produced by the beta-oxidation.

Next the acetyl-CoA feeds into citric acid cycle. In this process, it is converted to CO2. The energy from Acetyl-CoA is however extracted in the process. Specifically, each round generates 3 molecules of NADH, 1 molecule of FADH2 and 1 molecule of GTP. Since there are 12 molecules of Acetyl CoA, there will be 12 rounds of citric acid cycle. Therefore, there will be 12*3 = 36 molecules of NADH, 12*1 = 12 molecules of FADH2, and 12*1 = 12 molecules of GTP produced in the process.

Compiling both beta-oxidation and TCA, there are a total of 11+36 = 37 molecules of NADH, 11 + 12 = 23 molecules of FADH2, and 12 molecules of GTP produced.

GTP can be converted to ATP. For each molecule of GTP, one molecule of ATP is produced. Therefore, a total of 12 ATP molecules will be produced.

FADH2 and NADH feed into the Electron Transport Chain (ETC) and thus power the ATP synthase to produce ATP. For each molecule of FADH2, 1.5 ATP molecules are produced, and for each molecule of NADH, 2.5 ATP molecules are produced.

Therefore, for 37 NADH molecules 37 * 2.5 = 92.5 ATP molecules are produced. And for 23 FADH2 molecules 23 * 1.5 = 34.5 ATP molecules are produced.

Hence, a total of 92.5 + 34.5 + 12 = 139 ATP molecules are produced by the metabolism of 24:0 fatty acid.

asdfa
Answer #2

21......................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................................or 24

source: scarepant
answered by: spangebab